OpenStudy (anonymous):

x^2 + 5x + 4 = 0 Describe the solution to the equation by just determining the discriminat.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

find out the discriminant

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I have no clue how to tho

ganeshie8 (ganeshie8):

for a quadratic \(\large ax^2+bx+c\), \(\large b^2-4ac\) is called the discriminant

ganeshie8 (ganeshie8):

\(\large x^2 + 5x + 4 = 0 \) \(\large a = ?\) \(\large b = ?\) \(\large c = ?\)

OpenStudy (anonymous):

A= 0 B = 5 C = 4 ?

ganeshie8 (ganeshie8):

good try, but A = 1 okay ?

OpenStudy (anonymous):

oh because x = 1 dur

ganeshie8 (ganeshie8):

\(\large x^2 + 5x + 4 = 0\) is same as \(\large 1x^2 + 5x + 4 = 0\)

ganeshie8 (ganeshie8):

so, \(\large a = 1\) \(\large b = 5\) \(\large c = 4\)

ganeshie8 (ganeshie8):

find out the discriminant

OpenStudy (anonymous):

b2−4ac

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

yes, evaluate

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

9

ganeshie8 (ganeshie8):

\(\large a = 1\) \(\large b = 5\) \(\large c = 4\) discriminant = \(\large b^2 - 4ac = 5^2 - 4\times 1\times 4 = 25 - 16 = 9 \)

ganeshie8 (ganeshie8):

Yes ! discriminant = \(\large 9\)

OpenStudy (anonymous):

dang that was easy

ganeshie8 (ganeshie8):

yup :) but we're not done yet

ganeshie8 (ganeshie8):

we need to describe the type of solution by using this discriminant value

OpenStudy (anonymous):

oh

ganeshie8 (ganeshie8):

use below : 1) if discriminant > 0, then the equation will have "two different real solutions" 2) if discriminant = 0, then the equation will have "two same real solutions" 3) if discriminant < 0, then the equation will have "two complex solutions"

ganeshie8 (ganeshie8):

based on above conditions, wat do u think about the solutions for the given quadratic ?

OpenStudy (anonymous):

hmm do i plug in the 9 for x?

ganeshie8 (ganeshie8):

No, its much simpler than that

OpenStudy (anonymous):

ok please explain to me how because what you types up there made me totally lost

ganeshie8 (ganeshie8):

Since the discriminant = 9, which is greater than 0, the given equation will have "two different real solutions"

ganeshie8 (ganeshie8):

Moreover, since 9 is a perfect square, the solutions are rational as well.

OpenStudy (anonymous):

ok so that would be 9 < x < ?

ganeshie8 (ganeshie8):

Overall, the given equation will have "two different rational solutions"

ganeshie8 (ganeshie8):

we're done.

ganeshie8 (ganeshie8):

the question just asked u to describe the types of solutions

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok that works thank you i still am wondering tho about that 9 and you said its a perfect square so

ganeshie8 (ganeshie8):

4, 9, 16, 25.... are perfect squares... by any chance, do u remember the quadratic formula ?

OpenStudy (anonymous):

um is it like this\[b^2 +- √ \]

ganeshie8 (ganeshie8):

yes ! solutions = \(\large \dfrac{-b \pm \sqrt{\text{discriminant}}}{2a}\)

ganeshie8 (ganeshie8):

^^ discriminant goes and sits there

ganeshie8 (ganeshie8):

so, if discriminant is a perfect square, then there will not be any radicals in the solutions.

OpenStudy (anonymous):

OH

ganeshie8 (ganeshie8):

that means, the solutions will be rational numbers

ganeshie8 (ganeshie8):

for the present problem, just rephrase below : Since the discriminant = 9, which is greater than 0, the given equation will have "two different real solutions" Moreover, since 9 is a perfect square, the solutions are rational as well. Overall, the given equation will have "two different rational solutions"

OpenStudy (anonymous):

So for the question i should say something like this: b2−4ac=52−4×1×4=25−16=9

OpenStudy (anonymous):

oh ok thank you the almighty GANESHINE

ganeshie8 (ganeshie8):

yes put that also :)